You are the winner of a $10 million lottery where you get to choose between (option 1) $5 million today or (option 2) $500,000 at the beginning of each year, starting today, for the next 20 years. If the annual interest rate (EAR) is 9%, which option should you pick and why? What if the interest rate (EAR) is 6%?
Answer #1

PV of all future payments using annuity due formula = (1+9%)*500000*(1-(1+9%)-20)/9% = 4,975,057.39

Hence option of 5 million is better.

For rate = 6%
PV of all future payments using annuity due formula = (1+6%)*500000*(1-(1+6%)-20)/6% = 6,079,058.25

Hence in this case option of future payments is better.

Please Discuss in case of Doubt

Best of Luck. God Bless
Please Rate Well